OpenStudy (anonymous):

If 3(m – 2) + 4 – m = 18, what is the value of m?

OpenStudy (anonymous):

3m - 6 + 4 - m =18 3m - 2 - m = 18 2m -2 = 18 2m = 20 m =10

OpenStudy (anonymous):

thanx